# Directional Sign Loss: A Topology-Preserving Loss Function that Approximates the Sign of Finite Differences

### v1

@@ Abstract
Preserving critical topological features in learned latent spaces is a fundamental challenge in representation learning, particularly for topology-sensitive data. This paper introduces directional sign loss (DSL), a novel loss function that approximates the number of mismatches in the signs of finite differences between corresponding elements of two arrays. By penalizing discrepancies in critical points between input and reconstructed data, DSL encourages autoencoders and other learnable compressors to retain the topological features of the original data. We present the mathematical formulation, complexity analysis, and practical implementation of DSL, comparing its behavior to its non-differentiable counterpart and to other topological measures. Experiments on one-, two-, and three-dimensional data show that combining DSL with traditional loss functions preserves topological features more effectively than traditional losses alone. Moreover, DSL serves as a differentiable, efficient proxy for common topology-based metrics, enabling its use in gradient-based optimization frameworks. Code for all experiments is available here.
@@ 1 Introduction
Learned representations are central to modern machine learning, serving as compressed yet useful descriptions of data. While representations are often optimized for specific tasks, reconstructions are equally important, not only for downstream machine learning tasks, but also for human interpretability and applications. Standard loss functions consider only point-wise similarity, but often fail to capture topological structures in data. This can lead to reconstructions that appear to be faithful but miss geometric or structural features. While topology-aware learning methods, such as persistent homology, offer approaches to capturing these features, they are typically computationally inefficient and not differentiable, preventing their incorporation into gradient-based learning frameworks with large datasets.
Topological features capture the innate structure of data. These features include connected components, loops, and critical points (maxima, minima, and saddle points) (Carlsson, 2009) and are often obtained using topological data analysis (TDA) tools such as persistence diagrams. These features can be used for analyzing and visualizing the data structure, specifically for data where topology plays a vital role. Distance measures such as the Wasserstein distance describe the dissimilarity between the topological features of two pieces of data (Edelsbrunner and Harer, 2010). These measures help researchers draw out relevant information while filtering out noise or irrelevant variations in the data and are useful in machine learning (ML) tasks such as classifying or clustering topology-rich data (Maljovec et al., 2013; Yan et al., 2021). Scientific simulations, including those in fluid dynamics, physics, and weather prediction, leverage topological features to identify phenomena such as vortices, changes in atomic formation, and pressure extrema (Agarwal et al., 2021; Ramamurthi et al., 2022). Similarly, medical imaging often relies on capturing and comparing topological structures, such as gene interactions and tumors (Singh et al., 2023). Economists also examine financial networks through persistent features to detect market anomalies and gain insights into overall market behavior (Mojdehi et al., 2025; Eberle and Holder, 2007).
@@
  • Evaluations of learnable compressors, consisting of two simple autoencoders and one variational autoencoder, using this loss function, trained on 1-dimensional stock prices, 2-dimensional images, and 3-dimensional physics simulation states.
  • Theoretical and empirical evidence that this loss function can be used as a proxy for other established measures of topological dissimilarity.
In the rest of this paper, we introduce the background and related work (§ 2.2), describe the loss function and demonstrate its behavior on a simple example (§ 3), evaluate it by training several autoencoders and discussing the results (§ 4), and suggest future work and offer our conclusion (§ 5).
@@ 2.1.2 Topological data analysis
Critical points: A critical point of a continuous function $f$ is a point in the domain where $\nabla f=0$ (Edelsbrunner and Harer, 2010). However, in discrete settings such as sampled fields on a grid, the definition of critical points from continuous functions does not directly apply. For vector fields, identifying critical points can require detecting sign changes in the determinants of vectors within simplices formed from the vector field. This process involves evaluating whether replacing a vertex with zero leads to a determinant sign change, indicating a potential critical point (Xia et al., 2024). Because we focus on scalar fields (“tensors”), it is sufficient to approximate critical points by identifying locations where finite differences in some direction cross zero. If a critical point exists in an underlying continuous function $f$, then a zero crossing in the finite difference $\hat{f}(x+h)-\hat{f}(x)$ of the discrete approximation $\hat{f}$, using step size $h$, serves as an indicator of an approximate critical point with $O(h)$ error (Powell, 1981).
@@
Persistence diagrams: A persistence diagram is a topological summary that captures the birth and death of homological features in a filtered space. Given a filtration, a nested sequence of spaces defined by a function, topological features such as connected components, loops, and voids appear and disappear at specific function values. Each feature is represented as a tuple in the persistence diagram, where the first coordinate corresponds to the function value at which the feature first appears (birth) and the second coordinate corresponds to where it disappears (death). These pairs are created using the Elder Rule, which states that when two features merge, the older one persists while the newer one disappears. Since changes in topology occur at critical points of the filtration function, persistence diagrams implicitly encode information about critical points. Persistence diagrams provide a compact representation of topological structure. For a worked-through example, see Appendix 5.1.1.
@@
Distance measures: Several metrics quantify the dissimilarity between persistence diagrams, which in turn reflect differences in topological structure. Our work uses these metrics to evaluate the preservation of topological features between original and reconstructed data. The Wasserstein distance (Malinovskii, 1975) describes the cost of transforming one probability distribution into another, typically using norm order $p=1$ or $p=2$. The bottleneck distance is a special case of Wasserstein distance where $p=\infty$, capturing the largest discrepancy between the most dissimilar birth-death pairs. For persistence diagrams, these metrics compare differences between corresponding birth-death pairs in the two diagrams, implicitly reflecting variations in the underlying critical points of the filtration function. For a detailed description, see Appendix 5.1.3.
@@ 2.2 Related work
Lossy data compressors such as Di and Cappello (2016); Lindstrom (2014); Lakshminarasimhan et al. (2013); Lindstrom and Isenburg (2006) are usually developed to achieve a high compression ratio with an admissible fidelity loss. Other lossy compressors can perform certain operations in the compressed space (Agarwal et al., 2024a, b), some of which are differentiable (Agarwal et al., 2023). Some lossy compressors aim to guarantee the preservation of a controllable set of topological features during compression (Soler et al., 2018b; Yan et al., 2024; Xia et al., 2024). These compressors mainly focus on maintaining point-wise error along with high compression speed and ratio. While some explicitly target topological feature preservation, all of these compressors are “frozen” and are not designed to adapt to specific data distributions as learnable compressors are.
Autoencoders (AEs) have been widely used as learnable compression models. Traditional AEs and their common variants focus primarily on reconstructing the input while implicitly learning a latent space using a standard reconstruction loss, without explicitly considering the structural properties of the data.
To address this limitation, several studies have explored adding topological constraints into neural network training. Topological autoencoders (Moor et al., 2021) aim to preserve topological features of the input space in the latent representation. Persistence-weighted kernels (Kusano et al., 2016) and persistence landscapes (Bubenik, 2015) have been used to regularize learned representations and improve their structural retention.
Topological layers have been used to integrate persistent homology into learning objectives (Brüel-Gabrielsson et al., 2020) using level set and edge-based filters.
Topology-aware representational similarity analysis (tRSA) has appeared to characterize the topology of learned representations (Lin and Kriegeskorte, 2024). Such analysis techniques can be robust to noise and individual variability, suggesting potential benefits of preserving topology.
Our approach differs from existing methods in that it does not impose constraints on the latent space or require modifications to the architecture of the AE. Instead, our loss function directly encourages the preservation of topological features across compression and decompression. By avoiding explicit regularization of the latent representation, our method allows the AE to learn an efficient encoding while optimizing for structural fidelity in the reconstruction. This flexibility makes our approach applicable where efficiency and integrity must be balanced.
@@ 3.1 Algorithm
One way to capture the differences in critical points between two arrays of numbers (“tensors”) is to take their finite differences in all directions. If tensor $X\in\mathbb{R}^{n}$ were 1-dimensional, this would amount to computing $X_{2:n}-X_{1:n-1}$, and doing the same to the other tensor, say $Y$. The signs of these finite differences can be expressed as integers using $\mathrm{sign}(x)=\{1\;\mathrm{if}\;x>0,0\;\mathrm{if}\;x=0,-1\;\mathrm{if}\;x<0$. Then, the number of mismatches between the signs of their finite differences would be aggregated. This serves as a measure of the dissimilarity of their critical points as defined above, whether they be minima, maxima, or saddle points. In contrast to other metrics, this focuses only on
@@
the presence and location of the critical points and not on the values that construct them. However, counting mismatches directly would not be differentiable. Thus, we use a smooth function $\tanh$, with a sharpness parameter $s$, to approximate the $\mathrm{sign}$ function. Our experiments include comparisons between the approximate and exact metrics.
DSL computes the total mismatch in directional trends between tensors $X$ and $Y\in\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}$, which must have the same shape. One could be a reconstruction or approximation of the other.
For each dimension $i$ of the tensors from $1$ to $d$, the following steps are performed.
@@
    Approximate directional signs: The finite differences are scaled by a sharpness factor $s$, and the $\tanh$ function is applied, with larger values of $s$ producing a closer approximation to the $\mathrm{sign}$ function. Larger $s$ helps distinguish DSL from other loss functions such as MAE loss, but risks vanishing gradients, making the choice of $s$ a hyperparameter. Figure 1 shows $\tanh(sx)$ with different values of $s$. The user may use other $\mathrm{sign}$-like functions in DSL, such as $\mathrm{softsign}$ (Glorot and Bengio, 2010).
@@
The per-dimension mismatches are optionally weighted using weights $W\in\mathbb{R}^{d}$, allowing the user to select or emphasize certain dimensions. This also enables DSL to be applied to tensor fields by setting the elements $W_{(d-m):d}$ to zero, where $m$ is the dimensionality of an element in the field. These weighted mismatches are then accumulated into a scalar loss value. Pseudocode is shown in Algorithm 1. In the pseudocode, $a:b$ denotes all indices from $a$ to $b$. (:):(:)( : ) denotes all indices in some dimension, and $(:)*a$ denotes all indices for a sequence of $a$ dimensions.
Input: Tensors $X$ and $Y$ of shape $(n_{1},n_{2},\dots,n_{d})$, sharpness $s$
@@
       Compute finite differences:;
$D_{X}$ $\leftarrow X_{(:)*(i{-}1),2{:}n_{i},(:)*(d{-}i)}$
$\quad-X_{(:)*(i{-}1),1{:}(n_{i}{-}1),(:)*(d{-}i)}$
$D_{Y}$ $\leftarrow Y_{(:)*(i{-}1),2{:}n_{i},(:)*(d{-}i)}$
$\quad-Y_{(:)*(i{-}1),1{:}(n_{i}{-}1),(:)*(d{-}i)}$
Find approximate sign mismatches:;
@@
With optional weights $W$, accumulate mismatches:;
@@ 3.2 Time and memory complexity
We consider the time complexity using an idealized parallel computing model, with the availability of infinite parallel processors. If $X$ and $Y$ are $d$-dimensional arrays shaped $(n_{1},n_{2},\dots,n_{d})$, we could create all $\Delta$ arrays for each dimension in constant time. Then, all $\Delta$, containing a total of $d\prod_{i=1}^{d}(n_{i}-1)$ elements, are summed. Ignoring the subtraction of $1$ for large $n_{i}$, this takes $O(\log_{2}(d\prod_{i=1}^{d}n_{i}))$ time. Taking the logarithm of the product, we have $O(\log_{2}d+\sum_{i=1}^{d}\log_{2}n_{i})$. However this would take $O(d\prod_{i=1}^{d}n_{i})$ memory. Our implementation resembles the pseudocode in Algorithm 1, where the mismatches are summed separately per dimension. This has a complexity of
@@
For comparison, MSE has a time complexity of $O(\sum_{i=1}^{d}\log_{2}n_{i})$ and the same memory complexity. Making persistence diagrams and taking the Wasserstein distance between them takes $O(\prod_{i=1}^{d}n_{i}^{2})$ time and $O(\prod_{i=1}^{d}n_{i}^{d})$ memory.
Implementation notes:
  • In our implementation, we skip the main calculation for the first (batch) dimension and provide options for reduction (mean, sum, or none), resembling typical PyTorch loss functions.
  • If the option to match the units of the exact version of DSL (using $\mathrm{sign}$ instead of $\tanh$) is turned off, we scale the return value by the number of comparisons, making it easier to use in combination with conventional loss functions such as MSE loss.
  • As the choice of $s$ greatly affects the behavior of the loss function, we provide an approach to learn $s$ in Appendix 5.3.

### v3

@@ Abstract
Preserving topological features in learned latent spaces is a fundamental challenge in representation learning, particularly for topology-sensitive data. This paper introduces directional sign loss (DSL), an efficient, differentiable loss function that approximates the number of mismatches in the signs of finite differences between corresponding elements of two arrays. By penalizing discrepancies in critical points between input and reconstructed data, DSL encourages autoencoders and other learnable compressors to retain the topological features of the original data. We present the formulation and complexity analysis of DSL, comparing it to other non-differentiable topological measures. Experiments on multidimensional array data show that combining DSL with traditional loss functions preserves topological features more effectively than traditional losses alone. DSL serves as a differentiable, efficient proxy for common topology-based metrics, enabling topological feature preservation on previously impractical problem sizes and in a wider range of gradient-based optimization frameworks.
@@ 1 Introduction
Learned representations are central to modern machine learning, serving as compressed yet useful descriptions of data. While representations are often optimized for specific tasks, reconstructions are equally important, not only for downstream machine learning tasks, but also for human interpretability and applications. Standard loss functions consider only point-wise similarity and often fail to capture topological structures in data. This can lead to reconstructions that appear faithful but miss geometric or structural features. While topology-aware learning methods, such as persistent homology, offer approaches to capturing these features, they are typically computationally inefficient and not differentiable, preventing their incorporation into gradient-based learning frameworks with large datasets.
Topological features capture the innate structure of data. These features include connected components, loops, and critical points (maxima, minima, and saddle points) (Carlsson, 2009) and are often obtained using topological data analysis (TDA) tools such as persistence diagrams. These features can be used for analyzing and visualizing the data structure, specifically for data where topology plays a vital role. Distance measures such as the Wasserstein distance describe the dissimilarity between the topological features of two pieces of data (Edelsbrunner and Harer, 2010). These measures help researchers draw out relevant information while filtering out noise or irrelevant variations in the data and are useful in machine learning (ML) tasks such as classifying or clustering topology-rich data (Maljovec et al., 2013; Yan et al., 2021). Scientific simulations, including those in fluid dynamics, physics, and weather prediction, leverage topological features to identify phenomena such as vortices, changes in atomic formation, and pressure extrema (Agarwal et al., 2021; Ramamurthi et al., 2022). Similarly, medical imaging often relies on capturing and comparing topological structures, such as gene interactions and tumors (Singh et al., 2023). Economists also examine financial networks through persistent features to detect market anomalies and gain insights into overall market behavior (Mojdehi et al., 2025; Eberle and Holder, 2007). Preservation of topological invariants is often essential for reliable interpretation and downstream decision-making, even at the expense of numerical accuracy (Tierny, 2017).
@@
  • Evaluations of learnable compressors that show that this loss function can be used as a proxy for other measures of topological dissimilarity, consisting of two simple autoencoders and one variational autoencoder, using this loss function, trained on 1-dimensional arrays of stock prices, 2-dimensional images, and 3-dimensional physics simulation states.
In the rest of this paper, we introduce the background and related work (§ 2.2), describe the loss function and demonstrate its behavior on a simple example (§ 3), evaluate it by training several autoencoders and discussing the results (§ 4), discuss limitations (§ 5), and offer our conclusion (§ 6).
@@ 2.1.2 Topological data analysis
Critical points: A critical point of a continuous function $f$ is a point in the domain where $\nabla f=0$ (Edelsbrunner and Harer, 2010). However, in discrete settings such as sampled fields on a grid, the definition of critical points from continuous functions does not directly apply. For vector fields, identifying critical points can require detecting sign changes in the determinants of vectors within simplices formed from the vector field. This process involves evaluating whether replacing a vertex with zero leads to a determinant sign change, indicating a potential critical point (Xia et al., 2024). For scalar fields (“tensors”), it is sufficient to approximate critical points by identifying locations where finite differences in some direction cross zero. If a critical point exists in an underlying continuous function $f$, then a zero crossing in the finite difference $\hat{f}(x+h)-\hat{f}(x)$ of the discrete approximation $\hat{f}$, using step size $h$, serves as an indicator of an approximate critical point with $O(h)$ error (Powell, 1981).
@@
Persistence diagrams: A persistence diagram is a topological summary that captures the birth and death of homological features in a filtered space. Given a filtration, a nested sequence of spaces defined by a function, topological features such as connected components, loops, and voids appear and disappear at specific function values. Each feature is represented as a tuple in the persistence diagram, where the first coordinate corresponds to the function value at which the feature first appears (birth) and the second coordinate corresponds to where it disappears (death). When two features merge, the older one persists while the newer one disappears. Since changes in topology occur at critical points of the filtration function, persistence diagrams implicitly encode information about critical points. Persistence diagrams provide a compact representation of topological structure. For $d$-dimensional arrays, persistence diagrams can be constructed for homology orders up to ${d-1}$. However, the majority of persistent features in practice, particularly connected components (order 0), dominate the topological summary, with higher-order features becoming progressively less common in most real-world data. For a worked-through example, see Appendix 6.1.1.
@@
Distance measures: Several metrics quantify the dissimilarity between persistence diagrams, which in turn reflect differences in topological structure. Our work uses these metrics to evaluate the preservation of topological features between original and reconstructed data. The Wasserstein distance (Malinovskii, 1975) between persistence diagrams is the most common measure of topological dissimilarity. The Wasserstein distance describes the cost of transforming one probability distribution into another, typically using norm order $p=1$ or $p=2$. The bottleneck distance is a special case of Wasserstein distance where $p=\infty$, capturing the largest discrepancy between the most dissimilar birth-death pairs. For persistence diagrams, these metrics compare differences between corresponding birth-death pairs in the two diagrams, implicitly reflecting variations in the underlying critical points of the filtration function. For a detailed description, see Appendix 6.1.3. The functional distortion distance between merge trees measures both structural and functional differences by optimally matching nodes in the merge trees based on their scalar function values, and penalizing unmatched or structurally divergent nodes. The distance is computed by matching nodes from one tree to the closest nodes in the other (in terms of function value), summing the minimal differences, and adding a penalty for differences in the number of critical points. This metric directly compares the topological structure of scalar fields, rather than their persistence diagrams. The pairwise linear correlation distance quantifies the dissimilarity of the spatial arrangements of feature vectors in the original and reconstructed data. All pairwise Euclidean distances between feature vectors at each spatial location are computed. Then the correlation distance is one minus the Pearson correlation coefficient between the resulting distance matrices for the two datasets. This metric captures the preservation of the overall geometric arrangement of features.
@@ 2.2 Related work
Lossy data compressors such as Di and Cappello (2016); Lindstrom (2014); Lakshminarasimhan et al. (2013); Lindstrom and Isenburg (2006) are usually developed to achieve a high compression ratio with an admissible fidelity loss. Some lossy compressors prioritize other features, such as differentiability or compressed-space operations (Agarwal et al., 2023, 2024a, 2024b). Others aim to guarantee the preservation of a controllable set of topological features during compression (Soler et al., 2018b; Yan et al., 2024; Xia et al., 2024). These compressors mainly focus on maintaining point-wise error along with high compression speed and ratio. While some explicitly target topological feature preservation, all of these compressors are “frozen” and are not designed to adapt to specific data distributions as learnable compressors are.
Autoencoders (AEs) have been widely used as learnable compression models. Traditional AEs and their common variants focus primarily on reconstructing the input while learning a latent space using a standard reconstruction loss, without explicitly considering the structural properties of the data.
To address this limitation, several studies have explored adding topological constraints into neural network training. Topological autoencoders (Moor et al., 2021) aim to preserve topological features of the input space in the latent representation. Such autoencoders use measures such as representation topology divergence (RTD) (Trofimov et al., 2023) to describe topological dissimilarity between data manifolds and latent representations. While these methods are similar to our work, they are meant to capture the topological features of datasets rather than individual examples, relying on small samples (up to size $\approx 2^{8}$) of small examples ($\approx 2^{16}$ elements) to retain tractability due to their reliance on persistence diagrams or distance matrices.
Persistence-weighted kernels (Kusano et al., 2016) and persistence landscapes (Bubenik, 2015) have been used to regularize learned representations and improve their structural retention. Topological layers have been used to integrate persistent homology into learning objectives (Brüel-Gabrielsson et al., 2020) using level set and edge-based filters. Topology-aware representational similarity analysis (tRSA) has appeared to characterize the topology of learned representations (Lin and Kriegeskorte, 2024), showing robustness to noise and individual variability. Unlike these approaches, our method requires no explicit topological descriptors or architectural modifications, such as specialized layers.
Our approach differs from these methods in that it does not impose constraints on the latent space or require modifications to the architecture of the AE. While methods such as Moor et al. (2021); Trofimov et al. (2023) enforce topological alignment between data and latent spaces through divergence minimization, our loss function directly encourages the preservation of topological features across compression and decompression of entire inputs (and not samples thereof) without accessing the latent space. This flexibility makes our approach particularly applicable in scenarios where computational efficiency and topological integrity must be balanced.
@@ 3.1 Algorithm
One way to capture the differences in critical points between two arrays of numbers (“tensors”) is to take their finite differences in all directions. If tensor $X\in\mathbb{R}^{n}$ were one-dimensional, this would amount to computing $X_{2:n}-X_{1:n-1}$, and doing the same to the other tensor, say $Y$. The signs of these finite differences can be expressed as integers using $\mathrm{sign}(x)=\{1\;\mathrm{if}\;x>0,0\;\mathrm{if}\;x=0,-1\;\mathrm{if}\;x<0$. Then, the number of mismatches between the signs of their finite differences would be aggregated, serving as a measure of the dissimilarity of their critical points. In contrast to other metrics, this focuses only on the presence and location of the critical points and not on the values that construct them. However, counting mismatches directly would not be differentiable. Thus, we use a smooth function $\tanh$, with a sharpness parameter $s$, to approximate the $\mathrm{sign}$ function.
DSL computes the total mismatch in directional trends between tensors $X$ and $Y\in\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}$, which must have the same shape. One could be a reconstruction or approximation of the other. For each dimension $i$ of the tensors from 111 to $d$, the following steps are performed.
@@
    Approximate directional signs: The finite differences are scaled by a sharpness factor $s$, and the $\tanh$ function is applied, with larger values of $s$ producing a closer approximation to the $\mathrm{sign}$ function. Larger $s$ helps distinguish DSL from other loss functions such as MAE loss, but risks vanishing gradients, making the choice of $s$ a hyperparameter. Figure 1 shows $\tanh(sx)$ with different values of $s$. The user may use other $\mathrm{sign}$-like functions in DSL, such as $\mathrm{softsign}$ Glorot and Bengio (2010).
@@
[…]
Input: Tensors $X$ and $Y$ of shape $(n_{1},n_{2},\dots,n_{d})$, sharpness $s$, optional weights $W$
@@
   Compute finite differences:
$D_{X}\leftarrow X_{(:)*(i{-}1),2{:}n_{i},(:)*(d{-}i)}-X_{(:)*(i{-}1),1{:}(n_{i}{-}1),(:)*(d{-}i)}$
$D_{Y}\leftarrow Y_{(:)*(i{-}1),2{:}n_{i},(:)*(d{-}i)}-Y_{(:)*(i{-}1),1{:}(n_{i}{-}1),(:)*(d{-}i)}$
Find approximate sign mismatches:
@@
With optional weights $W$, accumulate mismatches:
@@ 3.2 Time and memory complexity
We consider the time complexity using an idealized parallel computing model, with the availability of infinite parallel processors to differentiate the parts that are inherently parallelizable and those that are sequential bottlenecks. Such assumptions are commonly employed in GPU and high-performance computing literature because they illustrate the theoretical scalability and potential limitations of parallel implementations. If $X$ and $Y$ are $d$-dimensional arrays shaped $(n_{1},n_{2},\dots,n_{d})$, then we can achieve a time complexity of $O(\log_{2}d+\sum_{i=1}^{d}\log_{2}n_{i})$, but this would take $O(d\prod_{i=1}^{d}n_{i})$ memory (See Appendix 6.2). Our implementation instead resembles the pseudocode in Algorithm 1, where the mismatches are summed separately per dimension. This has a complexity of
@@
For comparison, MSE has a time complexity of $O(\sum_{i=1}^{d}\log_{2}n_{i})$ and the same memory complexity. Making persistence diagrams and taking the Wasserstein distance between them takes $O(\prod_{i=1}^{d}n_{i}^{3})$ time and $O(\prod_{i=1}^{d}n_{i})$ memory. Merge tree distance computation between trees with $n$ and $m$ extrema has an optimized time complexity of $O(n^{2}m^{2}(n+m))$ using memoization, with $O(nm)$ memory required for storing intermediate matching costs (Morozov et al., 2013; Yan et al., 2022). Without memoization, this takes $O(2^{n+m})$ time. Pairwise linear correlation analysis between $n$ variables with $m$ samples each requires $O(n^{2}m)$ time for full matrix computation, reducible to $O(n^{2}\log m)$ using fast Fourier transform windowing techniques (Pearson, 1895). Memory complexity ranges from $O(n^{2})$ for storing all correlations to $O(1)$ when computing on-the-fly. Figure 2 shows experimental time and memory usage of these loss functions.
(a) time 1D
(b) memory 1D
(c) time 2D
(d) memory 2D
(e) legend
Figure 2: Time and peak memory usage from 1-dimensional and 2-dimensional array inputs using one AMD Ryzen 7 7800X3D CPU and one NVIDIA RTX 3090 GPU (24GB). Points are not shown if the time per example exceeds 1 second or GPU memory is exceeded. 2D inputs sizes show the number of elements per side of a square. Persistence diagram distance was the 2-order Wasserstein distance between order-0 (the cheapest) persistence diagrams. 3D inputs are shown in Appendix Figure 8.
Implementation notes: In our implementation, we skip the main calculation for the first (batch) dimension and provide options for reduction (mean, sum, or none), resembling typical PyTorch loss functions. If the option to match the units of the exact version of DSL (using $\mathrm{sign}$ instead of $\tanh$) is turned off, we scale the return value by the number of comparisons, making it easier to use in combination with conventional loss functions such as MSE loss. As the choice of $s$ greatly affects the behavior of the loss function, we provide an approach to learn $s$ in Appendix 6.5.